OpenStudy (anonymous):

MEDAL AND FAN! PLEASE HELP! how do i find the solution to 3Y = 15 + x x + 7Y = 5 ------------------------

jimthompson5910 (jim_thompson5910):

the first step is to isolate one of the variables let's say we pick on the second equation and solve for x x + 7y = 5 x+7y = 5 - 7y ... subtract 7y from both sides x = 5 - 7y so after solving `x + 7y = 5` for x, we end up with `x = 5 - 7y`

jimthompson5910 (jim_thompson5910):

now move onto the other equation we will replace every `x` with `5-7y` so... \[\Large 3y = 15 + x\] \[\Large 3y = 15 + \color{red}{x}\] \[\Large 3y = 15 + \color{red}{5-7y}\] \[\Large 3y = 15 + 5-7y\] do you know how to solve for y from here?

OpenStudy (anonymous):

no can you explain please

jimthompson5910 (jim_thompson5910):

does everything make sense so far up to this point?

OpenStudy (anonymous):

yes it makes sense. Once we get the X we substitute every X in the other equation with 5-7Y

jimthompson5910 (jim_thompson5910):

ok on the right side of the equation `3y = 15 + 5-7y` we can combine like terms, right?

OpenStudy (anonymous):

so that will make it 3y= 20-7y?

jimthompson5910 (jim_thompson5910):

yes, now we must move the y terms all to one side

OpenStudy (anonymous):

or will that make 10y = 20?

jimthompson5910 (jim_thompson5910):

good, you added 7y to both sides

jimthompson5910 (jim_thompson5910):

10y means 10 times y so 10 times ______ = 20 what must go in the blank? you can guess and check, or you can divide both sides by 10 to isolate y

OpenStudy (anonymous):

divided by 10 to both sides will give me Y = 2 ?

jimthompson5910 (jim_thompson5910):

yep y = 2 now use this to find x

jimthompson5910 (jim_thompson5910):

x = 5 - 7y x = 5 - 7*2 ... replace y with 2 x = ??

OpenStudy (anonymous):

-9 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the solution is (x,y) = (-9, 2) let's check that answer we do this by plugging it back into the original equations. It has to make ALL of the equations of the original system true

jimthompson5910 (jim_thompson5910):

3y = 15 + x 3y = 15 + (-9) ... replace x with -9 3*2 = 15 + (-9) ... replace y with 2 6 = 6 ... equation is true -------------------------------------- x + 7y = 5 -9 + 7y = 5 ... replace x with -9 -9 + 7*2 = 5 ... replace y with 2 -9 + 14 = 5 5 = 5... equation is true -------------------------------------- both original equations are true when (x,y) = (-9, 2) so the solution has been confirmed

jimthompson5910 (jim_thompson5910):

this check is highly recommended for any equation you solve because later on in algebra, you'll learn of cases where you'll get to an "answer" but it won't be the solution because it doesn't make the original equation true

OpenStudy (anonymous):

is this the only way to solve this equation? isn't there a way that goes like this. 3y = 15 + x -7y = -5 - x ------------- ? where you can cancel out x? sorry if this confuses you but i saw my teacher do this

jimthompson5910 (jim_thompson5910):

that's another way and that method is called the addition method or the elimination method

jimthompson5910 (jim_thompson5910):

notice on the right side of each equation, you can add them up and the +x and the -x combine to 0x, so the x terms go away completely

OpenStudy (anonymous):

can you show me the way to do it with this method? sorry if i'm bothering you. I have a test on this tomorrow

jimthompson5910 (jim_thompson5910):

it's no bother, so you're fine

jimthompson5910 (jim_thompson5910):

once you get it to this form 3y = 15 + x -7y = -5 - x you just add straight down

jimthompson5910 (jim_thompson5910):

3y plus -7y = -4y 15 plus -5 = 10 x plus -x = 0x

OpenStudy (anonymous):

-4y = 10 ?

jimthompson5910 (jim_thompson5910):

hmm something went wrong

jimthompson5910 (jim_thompson5910):

how did you get the -5? it should be +5 or just 5

jimthompson5910 (jim_thompson5910):

the system should be 3y = 15 + x -7y = 5 - x

jimthompson5910 (jim_thompson5910):

oh wait I'm misreading, let me try again

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

let's start with x + 7Y = 5 and subtract x from both sides, so we have just the y term on the left side so we now have this new system 3Y = 15 + x 7Y = 5 - x

jimthompson5910 (jim_thompson5910):

now add straight down

OpenStudy (anonymous):

10y = 20 but to get the bottom x to be a negative dont we have to times everything by -1?

jimthompson5910 (jim_thompson5910):

when you subtract x from both sides, it becomes negative on the right side

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):